OpenStudy (anonymous):

can you please help me with this problem? (-3, ___) is a solution of y = 4 - 3x.

OpenStudy (anonymous):

y = 4 - 3* (-3) = 4 +9 = 13 so, (-3,13) is a solution......

OpenStudy (anonymous):

ty